If the inverse demand curve of profit maximizing monopolist is given as P =1200 − 2Q , and cost function as

C = Q3 − 61.25Q2+1528.5Q + 2000, find equilibrium output level, monopolist price, and profit.





1
Expert's answer
2021-07-26T01:28:05-0400

Meaning

"p =1200Q \u2212 2Q ,"

"c = Q^3 \u2212 61.25Q^2+1528.5Q + 2000"

Now the inverse demand function is one where price has to be expressed as a function of quantity q

Now, we know that total revenue equals price times quantity, so that


"R(Q)= P(Q)\u00d7Q"

"P(Q)\u00d7Q=1200Q-2Q^2"


Dividing through by Q We get


"\frac{Q}{Q}P[Q]=\frac{1200Q}{Q}-\frac{2Q^2}{Q},"


"P[Q]=1200-2Q"


Which is the required inverse demand function



Let denote the profit function by. "\u03c0(Q)"

Now,

"\u03c0(Q)=R(Q)-TC(Q)"


"\u03c0(Q)=(1200Q-2Q^2)-(Q^3 \u2212 61.25Q^2+1528.5Q + 2000)"


"\u03c0(Q)=1200Q-2Q^2-Q^3 + 61.25Q^2-1528.5Q - 2000"


"\u03c0(Q)=-Q^3 + 59.25Q^2-328.5Q + 2000"

Which is the required profit function


Now finding profit maximizing output we need to find "Q^*" such that

"\u03c0(Q^*)=0....................(i)"

"\u03c0(Q^*)<0.................\n.(ii)"


Now

"\u03c0(Q)=-Q^3 + 59.25Q^2-328.5Q + 2000"

Differentiating both sides with respect to Q;

"\u03c0'(Q)=-3Q^2 + 118.5Q-328.5"


Differentiating both sides again with respect to Q;

"\u03c0''(Q)=-6Q + 118.5"

Now let the first satisfy condition (i)

"\u03c0(Q^*)=0....................(i)"

"-3(Q^*)^2s + 118.5Q^*-328.5=0"

"(Q^*)^2s - 39.5Q^*+109.5=0"

"2(Q^*)^2s -79Q^*+219=0"

"2(Q^*)^2s -6Q^*- 73Q^*+219=0"

"2Q^*(Q^*-3)-73(Q^*-3)=0"

"(2Q^*-73)(Q^*-3)=0"


Q= 3 or 36.5